Paul works for $6 an hour on weekdays, and makes $7 on weekends. Last week he made $250 for 40 hours. Determine how many hours he worked during the week and how many during the weekend.
Answer by checkley71(8403)

6X+7(40-X)=250
6X+280-7X=250
-X=250-280
-X=-30
X=30 HOURS WOEKED DURING THE WEEK FOR $6 PER HOUR
40-30=10 HOURS WORKED ON THE WEEKEND FOR $7 PER HOUR
PROOF 6*30+7*10=250
180+70=250
250=250